How high would a load of 32,000 cubic ft of grain fill a round silo 60 ft tall and 15 ft radius?
Answer by jim_thompson5910(35256)

Assuming this is a cylinder.


V = pi*r^2*h

32000 = pi*15^2*h

32000 = pi*225*h

32000 = 225pi*h

32000/(225pi) = h

1280/(9pi) = h

h = 1280/(9pi)

h = 45.270739368


The height of the load would be about 45.270739368 feet. This is a little over 3/4 of the way to the top.